# A Graph-Based Laser Path Solver Algorithm for Virtual Reality Laboratory Simulations

### v1

@@ 1 Introduction
femtoPro is an interactive laser laboratory simulation, enabling the safe conduct of complex experiments in an immersive virtual reality environment as we introduced elsewhere [1]. It features a realistic optical model that contains a mixture of geometrical and wave-optics characteristics to facilitate an accurate simulation in real time, ensuring a stable frame update rate of at least 90 frames per second (fps). Current methods for simulating light-wave propagation and interaction with matter use a variety of different approximations and models for a solution of Maxwell’s equations, with some also integrating quantum mechanical principles [2, 3, 4, 5]. Due to their significant computational demand, Maxwell solvers of high accuracy are typically used for designing optical experiments or simulating experimental data a-posteriori. In contrast, in a VR application, real-time calculation is essential, setting distinct requirements for femtoPro.
@@
Ray-tracing algorithms on GPU hardware have become highly efficient for simulating the behavior of light in real time, now even capable of handling computationally intensive phenomena like the dispersion of white light into color spectra and caustic effects [6]. Rather than tracing millions of rays in parallel, femtoPro simulates the propagation of one beam with a Gaussian intensity cross-section profile as it interacts with a sequence of optical elements (e.g., mirrors and lenses) in its pathway. A section of that path leading from one point of interaction (e.g., the laser source) to the next (e.g., a mirror) is defined as a beam segment. In order to implement femtosecond laser pulse properties, each beam segment contains complex-valued array structures that represent the frequency and time distributions of the electric field. This allows us to incorporate amplitude and phase information and handle interference phenomena. Given the specific needs to simulate these detailed laser pulse properties, we opted for a CPU-based algorithm over ray tracing on the GPU to efficiently manage associated calculations.
@@
The simulation is backed by a dynamic graph model that maps the formal network between beam segments of laser pulses propagating freely in space as edges, and optical elements (such as mirrors or lenses) that modify the laser pulse properties, as nodes. The graph’s topology evolves dynamically in real time, shaped by the underlying physical laws and interactions governing the system. Research into dynamic graph models in general has gained significant traction only recently, with a notable surge in interest observed during the COVID-19 pandemic [7]. The complexity of dynamic graph problems, potentially reaching NP-hardness, poses unique computational challenges and opens innovation avenues in multiple domains [8, 9]. In that context, we use an interdisciplinary approach, combining graph theory, optical physics, and practical optimization techniques to develop a solution tailored to the specific needs of an interactive real-time VR laser laboratory. By sharing insights into its conceptual framework, implementation, and performance evaluation, femtoPro not only showcases a distinctive application of dynamic graphs in real-time settings, but also highlights both challenges and solutions related to performance within such systems, serving as a reference for researchers and developers seeking to apply similar approaches in their work.
@@ 2 Requirements
This section outlines the features and requirements of femtoPro [10, 1] that form the basis for the underlying graph model and its implementation.
@@
A fundamental requirement of femtoPro is that experimental setups are fully interactive and simulated in real time, enabling immediate observation of the impacts of any interaction with the simulation. Similar to a real-world laboratory, there are various ways to manually adjust the position and/or orientation of an optical element via its opto-mechanical holder. Fine alignment is performed via adjustment screws which allow for vertical and horizontal tilt as well as displacement of an optical element. femtoPro’s goal to allow unrestricted exploration and experimentation through a freely-configurable experimental platform requires the capability of simulating any setup the users wish to construct, beyond just those predefined in literature. An experimental setup might be aligned in such way that a beam repeatedly reflects between two or more optical elements in a recurring cycle. In practice, particularly in optical cavities, the number of recurring reflections can approach infinity as the beam’s energy is decaying exponentially with increasing number of reflections [11]. Our model sets an arbitrary limit on the number of reflections that simplifies calculations but is not physically accurate. Figure 1 shows a finite bouncing scenario where each beam segment is annotated with an ascending number, indicating the unique and sequential path of laser light.
@@
The optical media previously described adhere to linear optics principles, where beams spatially overlapping at the media’s surface do not interact with each other, but instead preserve their individual properties. A nonlinear crystal (NLC) is a transmissive optical medium that also supports second-order nonlinear optical responses such as sum-frequency or second-harmonic generation. This means that two or more beams intersecting at the surface of an NLC may collaboratively create additional outgoing beams with frequencies different from those of the incident beams. It is important to note that while an outgoing beam might appear collinear and thus seem to be directly influenced by an incident beam, our model considers each transmitted beam to arise independently from the combined effects within the medium [1]. This approach offers a more precise representation of nonlinear light-matter interactions than other models that might assume direct interactions between overlapping incident beams.
@@
Figure 2 depicts a typical experimental setup for measuring nonlinear optical responses [11]. A laser beam is divided into two paths by a beam splitter. After being directed through several mirrors, the two beams are converged by a lens that aligns them to spatially overlap within an NLC. If the pulses also overlap temporally, their interaction with the NLC leads to the generation of a single sum-frequency beam via a second-order nonlinear response (purple arrow in the figure). In addition, each pulse on its own also creates a collinear second-harmonic beam. The relative time delay between the two split beams can be varied with nanometer precision using a motorized delay stage that controls the position of the attached mirrors (grey rectangle in the figure). For each time delay, the intensity of the generated beam is captured by a power meter, resulting in a second-order intensity autocorrelation curve that can be studied on a virtual laptop to determine pulse durations. The shape of the curve is influenced by the pulse shape and pulse duration. The complexity of such an experiment arises from the need to simulate the spatio-temporal nonlinear optical interactions of multiple interdependent beam segments, while simultaneously considering adjustments by the experimenter in real time.
@@ 3 Implementation
We implemented a dynamic graph model, which supports the modification of nodes as well as the addition and deletion of edges in real time. The graph’s entire topology is revealed only after all laser beams have traversed all relevant optical elements involved in forming their complete pathways through linear and nonlinear optical transformations. Algorithms for efficient graph traversal, such as breadth-first, depth-first, and A* search, are well-established in both literature and practice [12]. These methods, however, are generally tailored to static graphs where the entire topology is known in advance, making them unsuitable for a graph that is dynamically generated in real time. Furthermore, the conventional notion of “graph traversal” doesn’t precisely align with the concept of beams traversing through optical elements. In traditional graphs, edge generation and traversal are independent processes, with traversal typically taking place after the graph is completely constructed. However, in the context of a laser simulation, the generation and traversal of an edge are closely linked, consecutive steps within the ongoing graph generation process.
@@
Figure 5 depicts a simple example on how the order of processing nodes impacts the number of (unnecessary) calculations. Node two is a nonlinear optical element whose incident beams collectively impact the characteristics of any generated outgoing beam. This means that node two requires all available incident beams to perform a complete nonlinear transformation. In processing order variant A, however, node two is processed first. Since edge a and c are both required input for a nonlinear transformation, but the latter is not yet generated due to node three not yet being processed, the nonlinear transformation is based upon incomplete input and thus has to be performed twice. In contrast to this, variant B processed node three first, delivering all required input for the subsequent nonlinear transformation, therefore avoiding unnecessary and redundant calculations. To solve this problem, we implemented priority queuing [12] to ensure that all linear transformations are processed before any nonlinear transformations. This guarantees that a nonlinear optical element waits for complete input of linearly transformed beams before it performs a nonlinear transformation. However, if an experimental setup contains more than one nonlinear optical element and those elements depend on each other’s input of nonlinearly transformed beams, there is still a chance that a nonlinear transformation may be performed redundantly with incomplete input. This issue is not yet solved in femoPro, since experimental setups rarely require more than one nonlinear optical element. A potential solution could involve analyzing the pathways previously generated by linear transformations to determine an efficient processing order of multiple dependent nonlinear transformations.
@@
Efficient memory management techniques, although not included in the pseudocode, are crucial for the performance of the path solver’s implementation. A naive approach would be to allocate new memory for each beam segment that is newly generated by linear and nonlinear transformations within the node processing loop. Without deallocating the outdated beam segments from the previous frame, which have become irrelevant due to introduced changes, this would inevitably lead to memory overflow. To mitigate this issue, software frameworks such as game engines, including Unity, employ a garbage collection (GC) system that automatically frees up resources no longer in use by the application [13]. However, this process can temporarily halt other operations, leading to noticeable drops in the frame rate. These drops are particularly detrimental in a VR context because they can disrupt the immersive user experience, potentially causing severe discomfort such as motion sickness [14].
To ensure consistent and smooth performance, we needed a method to manage memory more efficiently than relying solely on GC, which led to the implementation of an extensive pooling system [15]. A pooling system preallocates a block of memory for objects that are frequently created and destroyed, in this case, the beam segments of the simulation. Instead of allocating new memory for each newly generated segment, segments are taken from this preallocated pool, and outdated segments from the previous frame are returned to the pool, making them available for reuse. When the pool reaches its capacity, it automatically expands by a predetermined size to accommodate more beam segments. It is important to note that pooling is helpful not only for managing beam segments but also for handling temporary data structures within physics calculations, especially those with a potentially large sample size. Initially, we utilized the Math.NET framework for data structures and operations [16], but it proved unsuitable as it allocated new memory for each computational operation rather than modifying variables in place. Since none of the solutions on the market met our specific needs for efficient real-time data processing, we developed our own pooled data structures that support in-place operations.
@@ 4.1 General performance analysis
We first perform a memory analysis of the path solver algorithm. An experimental setup can have a variable number of optical elements and beam segments. Each of those elements and segments has a fixed number of attributes to define their characteristics. Depending on its data type, storage of an attribute requires a certain amount of memory (in Bytes). As an example, a double-precision floating-point number requires 8 Bytes [17]. To determine the total memory cost of an optical element or beam segment, we sum up the type-specific cost of each of its attributes. Besides simple types such as integers and boolean flags, beam segments also have attributes of complex type, such as arrays. For instance, a beam’s frequency and wavelength values are stored in arrays with a specific sampling size. To determine the memory cost of those complex types, we simply have to multiply the per-element cost (e.g., 8 Bytes for double-precision floating-point numbers) with the sampling size. Conclusively, we calculate the memory requirements, $M$, of an experimental setup as follows
@@
Conservatively estimated, the total memory budget for applications on a mobile VR device is at least 1 GB [18]. From this available memory, each optical element requires approximately 4 KB, as calculated by inputting respective values into Equation 1b. Over 260,000 optical elements would be required to max out femtoPro’s available memory, a number that is far above the limit of several hundred optical elements imposed by physical and spatial constraints (see Section 2). For the currently implemented experimental objectives, a sampling size of $S=512$ is sufficient. As a result, by applying Equation 1c, we calculate that each beam segment requires approximately 16 KB of memory. Thereof, 98 % is allocated for array structures proportional to the sampling size, representing the largest chunk of pulse data.
@@
Runtime complexity is a fundamental concept in computer science that describes how the computational time requirements of an algorithm change as the size of the input data grows. It provides a way to quantify the efficiency of algorithms and predict their performance under different conditions. The notation used to express runtime complexity includes “Big Omega” ($\Omega$), which represents the lower bound or best-case scenario, and “Big O” ($O$), which denotes the upper bound or worst-case scenario [19].
@@
The overall runtime complexity of an algorithm is determined by the fastest-growing term. In the worst-case scenario for our simulation, the quadratic term from the nonlinear transformation, $C_{\text{nonlin}}$, dominates, resulting in a complexity of $O(b^{2})$. However, this quadratic worst-case complexity oversimplifies the nuanced impact of the distribution of incident beams $b$ across multiple nonlinear optical elements. While, for example, the complexity for one NLC with five incoming beams is quadratic ($255^{2}=$ 25), it is only linear for five NLCs with each having one incoming beam $\left((5)(1^{2})=5\right)$. Hence, the runtime complexity can be significantly less than $O(b^{2})$, depending on the beam distribution among nonlinear elements. Nonetheless, the presence of quadratic complexity in the worst-case scenario highlights the challenges inherent to nonlinear transformations.
@@
In this context, Table 1 presents the execution times per entity. All execution time measurements of Section 4.1 and 4.2 were gathered using the release version 0.7.0 of femtoPro for Android on the Pico 3 VR headset [20, 21]. We employed the stopwatch feature of C# [22] to mark specific parts of the simulation code for time measurements. For each term listed in the first column, we calculated the average execution time across 1,000 runs to minimize variations caused by other system components competing for CPU time (see column two). Notably, the execution time for $C_{\text{pair}}$, at approximately 113 µs, significantly exceeds other per-entity costs, which all remain below 60 µs.

### v2

@@ 1 Introduction
femtoPro is an interactive laser laboratory simulation, enabling the safe conduct of complex experiments in an immersive virtual reality environment as we introduced elsewhere [femtoProMain2023]. It features a realistic optical model that contains a mixture of geometrical and wave-optics characteristics to facilitate an accurate simulation in real time, ensuring a stable frame update rate of at least 90 frames per second (fps). A detailed discussion on the physical realism and accuracy of the underlying optical models is provided in a separate publication [femtoProSimulation2025]. Current methods for simulating light-wave propagation and interaction with matter use a variety of different approximations and models for a solution of Maxwell’s equations, with some also integrating quantum mechanical principles [SIM2009, QUANTUM, LUMERICAL, Meep2010]. Due to their significant computational demand, Maxwell solvers of high accuracy are typically used for designing optical experiments or simulating experimental data a-posteriori. In contrast, in a VR application, real-time calculation is essential, setting distinct requirements for femtoPro.
@@
Ray-tracing algorithms on GPU hardware have become highly efficient for simulating the behavior of light in real time, now even capable of handling computationally intensive phenomena like the dispersion of white light into color spectra and caustic effects [Raytracing2021]. Rather than tracing millions of rays in parallel, femtoPro simulates the propagation of one beam with a Gaussian intensity cross-section profile as it interacts with a sequence of optical elements (e.g., mirrors and lenses) in its pathway. A section of that path leading from one point of interaction (e.g., the laser source) to the next (e.g., a mirror) is defined as a beam segment. In order to implement femtosecond laser pulse properties, each beam segment contains complex-valued array structures that represent the frequency and time distributions of the electric field. This allows us to incorporate amplitude and phase information and handle interference phenomena. Given the specific needs to simulate these detailed laser pulse properties, we opted for a CPU-based algorithm over ray tracing on the GPU to efficiently manage associated calculations. While GPUs are optimized for highly parallel, uniform workloads, they are less suited for tasks involving tightly coupled data and dynamic structural changes. femtoPro’s evolving beam graph and interdependent field calculations require flexible memory handling and sequential control, making CPUs better suited to this class of simulation.
@@
The simulation is backed by a dynamic graph model that maps the formal network between beam segments of laser pulses propagating freely in space as edges, and optical elements (such as mirrors or lenses) that modify the laser pulse properties, as nodes. The graph’s topology evolves dynamically in real time, shaped by the underlying physical laws and interactions governing the system. Research into dynamic graph models in general has gained significant traction only recently, with a notable surge in interest observed during the COVID-19 pandemic [COVIDGRAPHS]. The complexity of dynamic graph problems, potentially reaching NP-hardness in general cases, poses unique computational challenges and opens innovation avenues in multiple domains [DynamicGraphHardness, ERLEBACH20211]. However, the specific problem addressed in femtoPro does not involve NP-hard complexity, as our current implementation follows a structured and constrained traversal approach tailored to real-time optical simulation. Future extensions may explore traversal optimizations that resemble NP-hard problems such as the Traveling Salesman Problem, for example to minimize redundant computations in complex nonlinear configurations [travelingSalesmanProblem]. We use an interdisciplinary approach, combining graph theory, optical physics, and practical optimization techniques to develop a solution tailored to the specific needs of an interactive real-time VR laser laboratory. By sharing insights into its conceptual framework, implementation, and performance evaluation, femtoPro not only showcases a distinctive application of dynamic graphs in real-time settings, but also highlights both challenges and solutions related to performance within such systems, serving as a reference for researchers and developers seeking to apply similar approaches in their work.
@@ 2 Requirements
This section outlines the features and requirements of femtoPro [femtoProArchitecture2023, femtoProMain2023] that form the basis for the underlying graph model and its implementation.
@@
A fundamental requirement of femtoPro is that experimental setups are fully interactive and simulated in real time, enabling immediate observation of the impacts of any interaction with the simulation. Similar to a real-world laboratory, there are various ways to manually adjust the position and/or orientation of an optical element via its opto-mechanical holder. Fine alignment is performed via adjustment screws which allow for vertical and horizontal tilt as well as displacement of an optical element.
For better understanding, Figure 1 shows a collection of in-game screenshots illustrating femtoPro’s core features in virtual reality as experienced by the user. The status of didactical tasks is presented on a whiteboard (Fig. 1a). The freely-configurable experimental setup (Fig. 1b) can be controlled and analyzed in various ways via a virtual laptop (Fig. 1c). The available experimental levels are listed in a level folder, which allows loading, saving, and restarting any given level (Fig. 1d). The virtual laptop, level folder, optical elements (e.g., Fig. 1e–i), and measurement devices (e.g., Fig. 1i) can be freely grabbed and placed on the experiment table. Figure 1j–n depicts further user interactions, such as navigating the lab environment via teleport (Fig. 1j), inputting numerical values via a numpad interface (Fig. 1k), locking an optical element onto the table via screws (Fig. 1l) as well as fine (Fig. 1m) and coarse (Fig. 1n) alignment of an optical element. A more detailed explanation of these features can be found elsewhere [femtoProArchitecture2023].
As illustrated in Figure 1, femtoPro offers a highly flexible, scalable, configurable, and freely interactable experimental environment. Users are not limited to predefined setups with fixed component positions; instead, they can assemble and precisely align any experiments with given optical elements, while still obtaining accurate results through our adaptable real-time physics simulation. The physical accuracy of the underlying optical models has been discussed and validated in a separate publication [femtoProSimulation2025].
Figure 1: Collection of screenshots showing femtoPro’s major features and interaction elements: (a) didactical tasks on the whiteboard, (b) experimental setup, (c) virtual laptop for controlling and analyzing experiments, (d) level folder for loading, saving, and restarting experimental levels, (e)–(i) optical elements, such as (e) laser source, (f) lens, (g) mirror, (h) beam blocker, (i) power meter, (j)–(n) user interactions, such as (j) teleporting, (k) inputting a number, (l) locking a screw, (m) fine and (n) coarse adjustment of an optical element. For snippets (e)–(n), a white background was added to enhance visual clarity. Further details are provided elsewhere [femtoProArchitecture2023].
femtoPro’s goal to allow unrestricted exploration and experimentation through a freely-configurable experimental platform requires the capability of simulating any setup the users wish to construct, beyond just those predefined in literature. An experimental setup might be aligned in such way that a beam repeatedly reflects between two or more optical elements in a recurring cycle. In practice, particularly in optical cavities, the number of recurring reflections can approach infinity as the beam’s energy is decaying exponentially with increasing number of reflections [trager2012]. Our model sets an arbitrary limit on the number of reflections that simplifies calculations but is not physically accurate. Figure 2 shows a finite bouncing scenario where each beam segment is annotated with an ascending number, indicating the unique and sequential path of laser light.
@@
The optical media previously described adhere to linear optics principles, where beams spatially overlapping at the media’s surface do not interact with each other, but instead preserve their individual properties. A nonlinear crystal (NLC) is a transmissive optical medium that also supports second-order nonlinear optical responses such as sum-frequency or second-harmonic generation. This means that two or more beams intersecting at the surface of an NLC may collaboratively create additional outgoing beams with frequencies different from those of the incident beams. It is important to note that while an outgoing beam might appear collinear and thus seem to be directly influenced by an incident beam, our model considers each transmitted beam to arise independently from the combined effects within the medium [femtoProMain2023]. This approach offers a more precise representation of nonlinear light-matter interactions than other models that might assume direct interactions between overlapping incident beams.
@@
Figure 3 depicts a typical experimental setup for measuring nonlinear optical responses [trager2012]. A laser beam is divided into two paths by a beam splitter. After being directed through several mirrors, the two beams are converged by a lens that aligns them to spatially overlap within an NLC. If the pulses also overlap temporally, their interaction with the NLC leads to the generation of a single sum-frequency beam via a second-order nonlinear response (purple arrow in the figure). In addition, each pulse on its own also creates a collinear second-harmonic beam. The relative time delay between the two split beams can be varied with nanometer precision using a motorized delay stage that controls the position of the attached mirrors (grey rectangle in the figure). For each time delay, the intensity of the generated beam is captured by a power meter, resulting in a second-order intensity autocorrelation curve that can be studied on a virtual laptop to determine pulse durations. The shape of the curve is influenced by the pulse shape and pulse duration. The complexity of such an experiment arises from the need to simulate the spatio-temporal nonlinear optical interactions of multiple interdependent beam segments, while simultaneously considering adjustments by the experimenter in real time.
@@ 3 Implementation
We implemented a dynamic graph model, which supports the modification of nodes as well as the addition and deletion of edges in real time. The graph’s entire topology is revealed only after all laser beams have traversed all relevant optical elements involved in forming their complete pathways through linear and nonlinear optical transformations. Algorithms for efficient graph traversal, such as breadth-first, depth-first, and A* search, are well-established in both literature and practice [GraphAlgosBook]. These methods, however, are generally tailored to static graphs where the entire topology is known in advance, making them unsuitable for a graph that is dynamically generated in real time. Furthermore, the conventional notion of “graph traversal” doesn’t precisely align with the concept of beams traversing through optical elements. In traditional graphs, edge generation and traversal are independent processes, with traversal typically taking place after the graph is completely constructed. However, in the context of a laser simulation, the generation and traversal of an edge are closely linked, consecutive steps within the ongoing graph generation process.
@@
Figure 6 depicts a simple example on how the order of processing nodes impacts the number of (unnecessary) calculations. Node two is a nonlinear optical element whose incident beams collectively impact the characteristics of any generated outgoing beam. This means that node two requires all available incident beams to perform a complete nonlinear transformation. In processing order variant A, however, node two is processed first. Since edge a and c are both required input for a nonlinear transformation, but the latter is not yet generated due to node three not yet being processed, the nonlinear transformation is based upon incomplete input and thus has to be performed twice. In contrast to this, variant B processed node three first, delivering all required input for the subsequent nonlinear transformation, therefore avoiding unnecessary and redundant calculations. To solve this problem, we implemented priority queuing [GraphAlgosBook] to ensure that all linear transformations are processed before any nonlinear transformations. This guarantees that a nonlinear optical element waits for complete input of linearly transformed beams before it performs a nonlinear transformation. However, if an experimental setup contains more than one nonlinear optical element and those elements depend on each other’s input of nonlinearly transformed beams, there is still a chance that a nonlinear transformation may be performed redundantly with incomplete input. This issue is not yet solved in femoPro, since experimental setups rarely require more than one nonlinear optical element. A potential solution could involve analyzing the pathways previously generated by linear transformations to determine an efficient processing order of multiple dependent nonlinear transformations.
@@
Efficient memory management techniques, although not included in the pseudocode, are crucial for the performance of the path solver’s implementation. A naive approach would be to allocate new memory for each beam segment that is newly generated by linear and nonlinear transformations within the node processing loop. Without deallocating the outdated beam segments from the previous frame, which have become irrelevant due to introduced changes, this would inevitably lead to memory overflow. To mitigate this issue, software frameworks such as game engines, including Unity, employ a garbage collection (GC) system that automatically frees up resources no longer in use by the application [gcInUnity2024]. However, this process can temporarily halt other operations, leading to noticeable drops in the frame rate. These drops are particularly detrimental in a VR context because they can disrupt the immersive user experience, potentially causing severe discomfort such as motion sickness [effectOfFramerate2023].
To ensure consistent and smooth performance, we needed a method to manage memory more efficiently than relying solely on GC, which led to the implementation of an extensive pooling system [Gregory2009]. A pooling system preallocates a block of memory for objects that are frequently created and destroyed, in this case, the beam segments of the simulation. Instead of allocating new memory for each newly generated segment, segments are taken from this preallocated pool, and outdated segments from the previous frame are returned to the pool, making them available for reuse. When the pool reaches its capacity, it automatically expands by a predetermined size to accommodate more beam segments. It is important to note that pooling is helpful not only for managing beam segments but also for handling temporary data structures within physics calculations, especially those with a potentially large sample size. Initially, we utilized the Math.NET framework for data structures and operations [mathNetWebsite], but it proved unsuitable as it allocated new memory for each computational operation rather than modifying variables in place. Since none of the solutions on the market met our specific needs for efficient real-time data processing, we developed our own pooled data structures that support in-place operations.
@@ 4.1 General performance analysis
We first perform a memory analysis of the path solver algorithm. An experimental setup can have a variable number of optical elements and beam segments. Each of those elements and segments has a fixed number of attributes to define their characteristics. Depending on its data type, storage of an attribute requires a certain amount of memory (in Bytes). As an example, a double-precision floating-point number requires 8 Bytes [sizeOfDouble]. To determine the total memory cost of an optical element or beam segment, we sum up the type-specific cost of each of its attributes. Besides simple types such as integers and boolean flags, beam segments also have attributes of complex type, such as arrays. For instance, a beam’s frequency and wavelength values are stored in arrays with a specific sampling size. To determine the memory cost of those complex types, we simply have to multiply the per-element cost (e.g., 8 Bytes for double-precision floating-point numbers) with the sampling size. Conclusively, we calculate the memory requirements, $M$, of an experimental setup as follows
@@
Conservatively estimated, the total memory budget for applications on a mobile VR device is at least 1 GB [metaQuestRamArticle]. From this available memory, each optical element requires approximately 4 KB, as calculated by inputting respective values into Equation 1b. Over 260,000 optical elements would be required to max out femtoPro’s available memory, a number that is far above the limit of several hundred optical elements imposed by physical and spatial constraints (see Section 2). For the currently implemented experimental objectives, a sampling size of $S=512$ is sufficient. As a result, by applying Equation 1c, we calculate that each beam segment requires approximately 16 KB of memory. Thereof, 98 % is allocated for array structures proportional to the sampling size, representing the largest chunk of pulse data.
@@
Runtime complexity is a fundamental concept in computer science that describes how the computational time requirements of an algorithm change as the size of the input data grows. It provides a way to quantify the efficiency of algorithms and predict their performance under different conditions. The notation used to express runtime complexity includes “Big Omega” ($\Omega$), which represents the lower bound or best-case scenario, and “Big O” ($O$), which denotes the upper bound or worst-case scenario [introToComplexity2022].
@@
The overall runtime complexity of an algorithm is determined by the fastest-growing term. In the worst-case scenario for our simulation, the quadratic term from the nonlinear transformation, $C_{\text{nonlin}}$, dominates, resulting in a complexity of $O(b^{2})$. However, this quadratic worst-case complexity oversimplifies the nuanced impact of the distribution of incident beams $b$ across multiple nonlinear optical elements. While, for example, the complexity for one NLC with five incoming beams is quadratic ($5^{2}=25$), it is only linear for five NLCs with each having one incoming beam $\left((5)(1^{2})=5\right)$. Hence, the runtime complexity can be significantly less than $O(b^{2})$, depending on the beam distribution among nonlinear elements. Nonetheless, the presence of quadratic complexity in the worst-case scenario highlights the challenges inherent to nonlinear transformations.
@@
In this context, Table 1 presents the execution times per entity. All execution time measurements of Section 4.1 and 4.2 were gathered using the release version 0.7.0 of femtoPro for Android on the Pico 3 VR headset [femtoProWebsite, pico3Website]. We employed the stopwatch feature of C# [stopwatchClass] to mark specific parts of the simulation code for time measurements. For each term listed in the first column, we calculated the average execution time across 1,000 runs to minimize variations caused by other system components competing for CPU time (see column two). Notably, the execution time for $C_{\text{pair}}$, at approximately 113 µs, significantly exceeds other per-entity costs, which all remain below 60 µs.